A remote-controlled boat can travel at 4 feet per second in calm water. For a specific trip, the boat traveled 48 feet with the current, and then it traveled 16 feet against the current in the same amount of time. How fast is the current?
Found 2 solutions by Boreal, rothauserc:
Answer by Boreal(15235)

distance=rate*time; rate= speed of boat (b)+ speed of current (c) or 4+c down stream and 4-c upstream
48=(4+c)* time
16=(4-c)*time
time is the same
time=48/(4+c)
time=16/(4-c)
set the two equal.
48/(4+c)=16/(4-c)
cross-multiply
192-48c=64+16c
64c=128
c=2 feet per second ANSWER
boat goes 6 ft/sec downstream and takes 8 seconds to go 48 feet
boat goes 2 ft/sec upstream and takes 8 seconds to go 16 feet

Answer by rothauserc(4718)

let x be the speed of the current and t be the time
:
1) (4+x) * t = 48
:
2) (4-x) * t = 16
:
solve equation 1 and equation 2 for t
:
t = 48/(4+x)
:
t = 16/(4-x)
:
set both equations = to each other
:
48/(4+x) = 16/(4-x)
:
cross multiply fractions
:
48 * (4-x) = 16 * (4+x)
:
3 * (4-x) = 4+x
:
12 - 3x = 4+x
:
4x = 8
:
x = 2
:
********************************
the current is 2 feet per second
********************************
: